OpenStudy (anonymous):

divide x^2-3x-14 / x-7

OpenStudy (lgbasallote):

interestingly...if the question was actually x^2 - 5x - 14 / x - 7 then it would be a lot easier

OpenStudy (anonymous):

May be they need quotient and remainder \[ \frac{x^2-3 x-14}{x-7}=4+x+\frac{14}{x-7} \]

OpenStudy (anonymous):

is that the answer?

Parth (parthkohli):

No, @KE8717504.

OpenStudy (anonymous):

so what's the answer?

Parth (parthkohli):

It's the one you will find when you attempt.